OpenStudy (princesssleelee):

Which function below is the inverse of f(x) = The quantity of four x minus three, over two.? f−1(x) = two over the quantity of four x minus three f−1(x) = two over the quantity of three minus four x f−1(x) = The quantity of two x minus three, over four f−1(x) = The quantity of two x plus three, over four

OpenStudy (anonymous):

b

OpenStudy (anonymous):

\[f(x)=\frac{4x-3}{2}\]?

OpenStudy (anonymous):

if so, it is definitely not B

OpenStudy (princesssleelee):

@satellite73 yes

OpenStudy (anonymous):

ok it is the same as solving say \[\frac{4x-3}{2}=7\]

OpenStudy (anonymous):

you know how to solve that?

OpenStudy (princesssleelee):

no lol

OpenStudy (anonymous):

oh then you are not going to like this !

OpenStudy (princesssleelee):

darn it

OpenStudy (anonymous):

\[\frac{4x-3}{2}=7\] multply by 2 get \[4x-3=14\]

OpenStudy (princesssleelee):

ok

OpenStudy (anonymous):

add \(3\) and get \[4x=17\] then divide by 4 to finish with \[x=\frac{17}{4}\]

OpenStudy (anonymous):

now we are going to do the same thing with variables you havef \[f(x)=\frac{4x-3}{2}\] lets write it as \[y=\frac{4x-3}{2}\]

OpenStudy (anonymous):

then switch x and y (because that is what the inverse does) and get \[x=\frac{4y-3}{2}\] solve for \(y\) the steps are the same as above

OpenStudy (anonymous):

\[\frac{4y-3}{2}=x\] multiply by 2 get \[4y-3=2x\]

OpenStudy (anonymous):

add 3 get \[4y=2x+3\] then divide by 4 to get \[y=\frac{2x+3}{4}\]

OpenStudy (princesssleelee):

very very confusing but ill try

OpenStudy (anonymous):

and that is your inverse \[f^{-1}(x)=\frac{2x+3}{4}\]

OpenStudy (anonymous):

it is confusing if you cannot solve \[\frac{4x-3}{2}=7\] but the steps are only three : multiply by 2 add 3 divide by 4

OpenStudy (anonymous):

that is what the inverse function \[f^{-1}(x)=\frac{2x+3}{4}\] does

OpenStudy (princesssleelee):

x=−32+2y

OpenStudy (princesssleelee):

i have no clue lol

OpenStudy (anonymous):

you lost me there i wrote the steps above, see if you can follow them if not let me know all is written out i did not skip any

OpenStudy (princesssleelee):

x=4.25

OpenStudy (princesssleelee):

@satellite73

OpenStudy (anonymous):

yes that is with the numbers now take a look at the same steps with i wrote it with variables

OpenStudy (princesssleelee):

i dont know what to do there

OpenStudy (anonymous):

i did it for you follow the steps to the answer

OpenStudy (princesssleelee):

do i plug in the x

OpenStudy (anonymous):

follow the post starting with "now we are going to do the same thing with variables"

OpenStudy (princesssleelee):

i have no idea what I do with the steps?

OpenStudy (anonymous):

follow them i did it for you

OpenStudy (princesssleelee):

i cant figure it out to save my life